f(x) = 5x/4 + 2/3
What is f(a) if f(a+3)= 5?
Please help me

Answer by rothauserc(4718)

f(x) = 5x/4 + 2/3
:
f(a+3) = 5 = (5 * (a+3) / 4) + 2/3
:
12 is our common denominator
:
3 * (5 * (a+3)) + 8 = 5 * 12 = 60
:
15a + 45 + 8 = 60
:
15a = 7
:
a = 7/15
:
*******************************************
f(a) = 5*(7/15) + 2/3 = 7/3 + 2/3 = 9/3 = 3
*******************************************
: